Two groups, A and B, with 10 people in each are in an experiment, and each person drops out of the experiment independently with probability 0.2. Find the probability at least 9 people complete the experiment in either A or B, but not both.
Answer #1

let C and D are event so that at least 9 person complete the experiment in A and B

P(C) =P(X=9)+P(X=10)=(10C9)*(0.8)9(0.2)1+(10C10)*(0.8)10(0.2)0 =0.3758

P(D) =0.3758

since event C and D are independent,

P(C n D) =P(C)*P(D) =0.3758*0.3758=0.1412

probability at least 9 people complete the experiment in either A or B, but not both

=P(C u D) =P(C)+P(D)-2*P(C n D) =0.3758+0.3758-2*0.1412=0.4692